airplane travels 5580 km against the wind in 9 hoirs and 6750 km with the wind in the same amount of time. what is the rate of the plane kn still air and what is the rate of the wind
Found 3 solutions by Boreal, Alan3354, ikleyn:
Answer by Boreal(15235)

p-w=5580/9=620 km/hr
p+w=6750/9=750 km/hr
add them
2p=1370
p=685 km/hr
substitute into first one
685-w=620
w=65 km/hr
check with second one
685+65=750
plane goes 620 km/hr in still air
the wind speed is 65 km/hr.

Answer by Alan3354(69443)

"plane in still air" is airspeed.

Answer by ikleyn(53763)

.
You can find similar fully solved similar problems on upstream and downstream round trips with detailed solutions in the lessons
    - Wind and Current problems
    - More problems on upstream and downstream round trips
    - Selected problems from the archive on the boat floating Upstream and Downstream
in this site.